# disposition

### v2

@@ 5.2 Scalarization results for the concept of weak efficiency
Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid cone, $\bar{x}\in\Omega$, and $A:=\mathbb{A}(\bar{x})=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is nontrivial. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively. Suppose $S_{-K}$ is solid and
$S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset.$ (5.1)
@@
Then, there is $(x^{*},\alpha)\in K^{a\circ}$ such that $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.14 is valid.
Proof 5.20.
@@
Under our assumptions, by Proposition 4.13 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that $A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$. Taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset,$ hence $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Since $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, Proposition 5.14 yields the rest of the conclusion.
Remark 5.21.
In view of Remark 5.18, the point $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ given in the conclusion of Theorem 5.19 belongs to ${\mathrm{W}Eff}(\Omega\mid f,K)$.
@@
In the convex case (i.e., $K$ and $\mathbb{A}(\bar{x})$ are convex cones), using Proposition 4.14 we are able to replace the separation condition (5.1) by $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, as shown in the following result.
Corollary 5.22 (convex case).
@@
Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid, convex cone, $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is nontrivial and convex. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively, and $S_{K}$ is solid.
Then, the conclusion of Theorem 5.19 holds.
Proof 5.23.
Follows from Theorem 5.19 and Proposition 4.14.
Remark 5.24.
@@
Recall that, for any set $\Omega\subseteq Y,$ the quasi-interior of $\Omega$ is the set ${\mathrm{q}i}\,\Omega:=\{x\in\Omega\mid\operatorname{cl}(\mathbb{R}_{+}\cdot(\Omega-x))=Y\}$. In order to ensure that the assumption $A$ is nontrivial and convex from Corollary 5.22 is satisfied it is sufficient to impose that $Y$ is a real Hausdorff locally convex space and $f$ is $K$-convexlike (i.e., $f[\Omega]+K$ is convex). Indeed, assume that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$. Since $f$ is $K$-convexlike, we infer that $A=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is a convex cone in $Y$. By [27, Lem. 3.3 (4)] we know that $f(\bar{x})\notin\operatorname{int}(f[\Omega]+K)$. Since the interior and the quasi interior of a solid, convex set $f[\Omega]+K$ in a real Hausdorff locally convex space coincide (see [23, p. 66-67]), we have $A\subseteq{\mathrm{c}l}(\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x})))\neq Y$. Moreover, since $K\neq\{0\}$ we have $A\neq\{0\}$ as well, hence $A$ is nontrivial.
@@ 5.3 Scalarization results for the concept of proper efficiency with respect to a cone-valued map
We are ready to present some further main scalarization results which ensure that an $\mathbb{A}$-properly efficient solution of the vector problem ($P$) is a solution of the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) for some parameters $a\in Y$ and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$.
Theorem 5.25.
@@
$({\mathrm{c}l}\,S_{A}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$ (5.2)
Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.12 is valid.
Proof 5.26.
By Proposition 4.15 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $A\cap-C_{\psi}(x^{*},\alpha)=\{0\}$ and $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$. Then, taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-C_{\psi}(x^{*},\alpha)\subseteq A\cap-C_{\psi}(x^{*},\alpha)=\{0\},$ hence $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Finally, observing that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
Remark 5.27.
In view of Remark 5.18, if $K$ is solid or $\psi$ is continuous, then the point $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ given in the conclusion of Theorem 5.25 belongs to ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
Using Proposition 4.1 we are able to replace the separation condition (5.2) by $0\notin\operatorname{cl}\,S_{-K}$ and $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$ as well as some more specific conditions that we formalize below.
Corollary 5.28.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}$, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$. Suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, and that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact. Moreover, assume that one of the following statements is valid:
@@
    There is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (4.4) is valid, and one of the conditions (4.5)–(4.8) holds.
Then, the conclusion of Theorem 5.25 is valid.
Proof 5.29.
Using Proposition 4.1 a) and b) respectively, under our standing assumptions a) and b), we get that $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$ and $0\notin\operatorname{cl}\,S_{-K}$ imply (5.2). Note that the nontrivialness of $A$ follows from $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$, $K$ is nontrivial and $A\cap(-K)=\{0\}$. Then, the result follows from Theorem 5.25.
In the remaining part of the section, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 4.11 and Lemma 4.9.
Corollary 5.30.
Assume that ($Y,||\cdot||$) is a real normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$. Moreover, assume that one of the following statements is valid:
@@
    $B_{A}(||\cdot||)$ and ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ are compact.
@@
    $A$ is weakly closed, and either $Y$ is reflexive or ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ and $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ are weakly compact and $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
@@
Then, there is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.12, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid.
Proof 5.31.
@@
Note that $\varphi_{x^{*},\alpha}$ from the conclusion is constructed with respect to the equivalent norm $||\cdot||^{\prime}$. Since $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ and $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$, using Proposition 4.11 a) and b) respectively, under our standing assumptions a) and b), one gets that $({\mathrm{c}l}\,S_{A}^{0}(||\cdot||^{\prime}))\cap({\mathrm{c}l}\,S_{-K}(||\cdot||^{\prime}))=\emptyset$ for some norm $||\cdot||^{\prime}\sim||\cdot||$. Applying Lemma 4.9 ($3^{\circ}$, $4^{\circ}$), it follows that ${\mathrm{c}l}\,S_{-K}(||\cdot||^{\prime})$ is compact under a), respectively, weakly compact under b). Consequently, the conclusion follows by applying Theorem 5.25 for $Y$ endowed with the norm $\psi(\cdot):=||\cdot||^{\prime}$ under a), respectively, endowed with the weak topology and $\psi(\cdot):=||\cdot||^{\prime}$ under b).
@@
For the finite dimensional case, we obtain from Corollary 5.30 the following result.
Corollary 5.32.
@@
Assume that ($Y,||\cdot||$) is a real finite dimensional normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, pointed, convex cone, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a closed cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
Then, the conclusion of Corollary 5.30 holds.
Proof 5.33.
@@
By the well-known Krein-Rutman theorem (see, e.g., [49, Th. 2.4.7]) we know that $K^{\#}\neq\emptyset$, and by [29, Th. 3.1 ($4^{\circ}$)] it follows that $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$. So, the assumptions of Corollary 5.30 are satisfied.
Remark 5.34.
@@
Besides the proper efficiency concepts in the senses of Benson, Borwein and Hurwicz, we gain also scalarization results for the proper efficiency concept in the sense of Henig taking into account the known relationships between proper efficiency concepts in vector optimization. In order to emphasize this fact, consider a nontrivial cone $K\subseteq Y$ with normlike-base $B_{K}(\psi)$ and suppose that either $K$ is solid or $\psi$ is continuous. Assume that ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)\subseteq{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$. Taking into account Remark 5.27, in this subsection, we have provided sufficient conditions for the equivalence of the following assertions:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
  • $3^{\circ}$.
@@
    There is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 5.12 is valid),
@@
    There is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 5.12, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid).
Remark 5.35.
Under the assumptions in Corollary 5.28 a) (in particular the convexity of $A$), we were able to replace the separation condition (5.2) by $0\notin{\mathrm{c}l}\,S_{-K}$ and $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$ (which implies the condition $A\cap(-K)=\{0\}$) in our scalarization results. In contrast, the scalarization result (for Henig and Benson proper efficiency concepts) by García-Castaño, Melguizo-Padial and Parzanese [16, Cor. 3.10] in real normed spaces still contains a separation condition (SSP), also under convexity of $A$ (ensured by a convexlikness assumption on $f$).

### v3

@@ 4.3 Scalarization results for the concept of weak efficiency
Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid cone, $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot))$, $\bar{x}\in\Omega$, and $A:=\mathbb{A}(\bar{x})$ is nontrivial. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively. Suppose $S_{-K}$ is solid and
$S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset.$ (23)
@@
Then, there is $(x^{*},\alpha)\in K^{a\circ}$ such that $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing, any of the equivalent assertions given in Proposition 18 is valid.
Proof.
@@
Under our assumptions, by Proposition 24 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that $A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$. Taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset,$ hence $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Since $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, Proposition 18 yields the rest of the conclusion. ∎
@@
In the convex case (i.e., $K$ and $\mathbb{A}(\bar{x})$ are convex cones), using Proposition 25 we are able to replace the separation condition (23) by $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$, as shown in the following result.
Corollary 30 (convex case).
@@
Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid, convex cone, $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot))$, $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is nontrivial and convex. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively, and $S_{K}$ is solid.
Then, the conclusion of Theorem 29 holds.
Proof.
Follows from Theorem 29 and Proposition 25. ∎
Remark 17.
In view of Remark 12, since the point $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ in the conclusion of Theorem 29 belongs to ${\rm WEff}(\Omega\mid f,K)$, the result provides sufficient conditions for weak efficiency. Under the assumptions of Corollary 30 (including convexity of both cones) and taking into account Remark 12, we derive the equivalence of the following assertions:
  • $1^{\circ}$
    $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$.
  • $2^{\circ}$
    There is $(x^{*},\alpha)\in K^{a\circ}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 17 is valid, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing).
From this perspective, $2^{\circ}$ is a necessary and sufficient condition for the weak efficiency of $\bar{x}\in\Omega$.
Remark 18.
@@
Recall that, for any set $\Omega\subseteq Y,$ the quasi-interior of $\Omega$ is the set ${\rm qi}\,\Omega:=\{x\in\Omega\mid\operatorname{cl}(\mathbb{R}_{+}\cdot(\Omega-x))=Y\}$. In order to ensure that the assumption $A$ is nontrivial and convex from Corollary 30 is satisfied it is sufficient to impose that $Y$ is a real Hausdorff locally convex space and $f$ is $K$-convexlike (i.e., $f[\Omega]+K$ is convex). Indeed, assume that $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$. Since $f$ is $K$-convexlike, we infer that $A=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is a convex cone in $Y$. By [52, Lem. 3.3 (4)] we know that $f(\bar{x})\notin\operatorname{int}(f[\Omega]+K)$. Since the interior and the quasi interior of a solid, convex set $f[\Omega]+K$ in a real Hausdorff locally convex space coincide (see [21, p. 66-67]), we have $A\subseteq{\rm cl}(\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x})))\neq Y$. Moreover, since $K\neq\{0\}$ we have $A\neq\{0\}$ as well, hence $A$ is nontrivial.
Remark 19.
In this and upcoming sections, we consider a seminorm $\psi$ which is positive on $K\setminus\{0\}$ and on $\mathbb{A}(\bar{x})\setminus\{0\}$. In Example 2(a,b) we discussed ways to construct such seminorms. Of course, changing the $\bar{x}$ may alter the cone $\mathbb{A}(\bar{x})$, so one may require another seminorm $\psi$ (at least if $\psi$ is not a norm). In general, one must be careful when choosing a seminorm. Ideally, one would want a seminorm $\psi$ that is positive on $K\setminus\{0\}$ and on $\mathbb{A}(\bar{x})\setminus\{0\}$ for all points $\bar{x}\in\Omega$ which are relevant in the results (e.g. $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$ in Corollary 30). This is because the scalarized problems appearing in the conclusions of the results are then defined with respect to the same seminorm $\psi$. Note that any norm has the desired positivity property since it is positive on $Y\setminus\{0\}$.
@@ 4.4 Scalarization results for the concept of proper efficiency with respect to a cone-valued map
We are ready to present some further main scalarization results which ensure that an $\mathbb{A}$-properly efficient solution of the vector problem ($P$) is a solution of the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) for some parameters $a\in Y$ and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$. Remember that $\mathbb{A}:\Omega\to 2^{Y}$ is a cone-valued map.
Theorem 31.
@@
$({\rm cl}\,S_{A}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$ (24)
Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing, any of the equivalent assertions given in Proposition 17 is valid.
Proof.
By Proposition 26 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $A\cap-C_{\psi}(x^{*},\alpha)=\{0\}$ and $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$. Then, taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-C_{\psi}(x^{*},\alpha)\subseteq A\cap-C_{\psi}(x^{*},\alpha)=\{0\},$ hence $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Finally, observing that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 1 ($7^{\circ}$), Proposition 17 yields the rest of the conclusion. ∎
Remark 20.
In view of Remark 12, if $K$ is solid or $\psi$ is continuous, then the point $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ in the conclusion of Theorem 31 belongs to ${\rm PEff}_{He}(\Omega\mid f,K)$, hence the result provides sufficient conditions for proper efficiency in the sense of Henig.
Using Proposition 19 we are able to replace the separation condition (24) by $0\notin\operatorname{cl}\,S_{-K}$ and $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$ as well as some more specific conditions that we formalize below.
Corollary 32.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}$, $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$. Suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, and that one of the sets ${\rm cl}\,S_{-K}$ and ${\rm cl}\,S_{A}^{0}$ is compact. Moreover, assume that one of the following statements is valid:
@@
    There is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (14) is valid, and one of the conditions (15)–(18) holds.
Then, the conclusion of Theorem 31 is valid.
Proof.
Using Proposition 19 a) and b) respectively, under our standing assumptions a) and b), we get that $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$ and $0\notin\operatorname{cl}\,S_{-K}$ imply (24). Note that the nontrivialness of $A$ follows from $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$, $K$ is nontrivial and $A\cap(-K)=\{0\}$. Then, the result follows from Theorem 31. ∎
In the remaining part of the section, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 23 and Lemma 22.
Corollary 33.
Assume that ($Y,||\cdot||$) is a real normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$, $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$. Moreover, assume that one of the following statements is valid:
@@
    $B_{A}(||\cdot||)$ and ${\rm cl}\,S_{-K}(||\cdot||)$ are compact.
@@
    $A$ is weakly closed, and either $Y$ is reflexive or ${\rm cl}\,S_{-K}(||\cdot||)$ and $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ are weakly compact and $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
@@
Then, there is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$. Moreover, if $\varphi$ is $-C_{||\cdot||^{\prime}}(x^{*},\alpha)$ representing, any of the equivalent assertions given in Proposition 17, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid.
Proof.
@@
Note that $\varphi_{x^{*},\alpha}$ from the conclusion is constructed with respect to the equivalent norm $||\cdot||^{\prime}$. Since $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ and $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$, using Proposition 23 a) and b) respectively, under our standing assumptions a) and b), one gets that $({\rm cl}\,S_{A}^{0}(||\cdot||^{\prime}))\cap({\rm cl}\,S_{-K}(||\cdot||^{\prime}))=\emptyset$ for some norm $||\cdot||^{\prime}\sim||\cdot||$. Applying Lemma 22 ($3^{\circ}$, $4^{\circ}$), it follows that ${\rm cl}\,S_{-K}(||\cdot||^{\prime})$ is compact under a), respectively, weakly compact under b). Consequently, the conclusion follows by applying Theorem 31 for $Y$ endowed with the norm $\psi(\cdot):=||\cdot||^{\prime}$ under a), respectively, endowed with the weak topology and $\psi(\cdot):=||\cdot||^{\prime}$ under b). ∎
@@
For the finite dimensional case, we obtain from Corollary 33 the following result.
Corollary 34.
@@
Assume that ($Y,||\cdot||$) is a real finite dimensional normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, pointed, convex cone, $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a closed cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
Then, the conclusion of Corollary 33 holds.
Proof.
@@
By the well-known Krein-Rutman theorem (see, e.g., [42, Th. 2.4.7]) we know that $K^{\#}\neq\emptyset$, and by [15, Th. 3.1 ($4^{\circ}$)] it follows that $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$. So, the assumptions of Corollary 33 are satisfied. ∎
Remark 21.
@@
Note that Corollaries 32, 33 and 34 provide necessary conditions for $\mathbb{A}$-proper efficiency (in the sense of Definition 7), including the proper efficiency concepts in the senses of Benson, Borwein and Hurwicz (as mentioned in Remark 5). Moreover, we gain scalarization results for the proper efficiency concept in the sense of Henig taking into account the known relationships between proper efficiency concepts in vector optimization. In order to emphasize this fact, consider a nontrivial cone $K\subseteq Y$ with normlike-base $B_{K}(\psi)$ and suppose that either $K$ is solid or $\psi$ is continuous. Assume that ${\rm PEff}_{He}(\Omega\mid f,K)\subseteq{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$. Taking into account Remark 20, in this subsection, we have provided sufficient conditions for the equivalence of the following assertions:
  • $1^{\circ}$
    $\bar{x}\in{\rm PEff}_{He}(\Omega\mid f,K)$.
  • $2^{\circ}$
    $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$.
  • $3^{\circ}$
@@
    There is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 17 is valid, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing),
@@
    There is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 17, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid, if $\varphi$ is $-C_{||\cdot||^{\prime}}(x^{*},\alpha)$ representing).
From this perspective, $3^{\circ}$ (respectively, $3^{\circ}$’) is a necessary and sufficient condition for the proper efficiency of $\bar{x}\in\Omega$.
Remark 22.
Under the assumptions in Corollary 32 a) (in particular the convexity of $A$), we were able to replace the separation condition (24) by $0\notin{\rm cl}\,S_{-K}$ and $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$ (which implies the condition $A\cap(-K)=\{0\}$) in our scalarization results. In contrast, the scalarization result (for Henig and Benson proper efficiency concepts) by García-Castaño, Melguizo-Padial and Parzanese [1, Cor. 3.10] in real normed spaces still contains a separation condition (SSP), also under convexity of $A$ (ensured by a convexlikness assumption on $f$).
It is worth mentioning that a novel aspect of our work in comparison to other conic scalarization related works in the normed space framework (e.g. [8, 9], [1]) is the treatment of equivalent norms in the scalarization results (as done in Corollaries 33 and 34).